Instrument Log 

SAS-ZE          C1                  CF Lamont, OK

Date

Technician

Comments

4/25/2016

MS

A DAQ window was up and wouldn't let me close it out.  Restarted program and RTS at 1333 GMT.  DAQ window did not show back up.

4/22/2016

MS

A DAQ window was up and would not let me close it out.  Restarted program and  RTS at 1333 GMT.  DAQ window never showed back up.

3/3/2016

MS

A pop up window had occurred causing instruments program to stop running at 0700 GMT.  Sent the mentor a screen shot of the pop up window.  At 1417 GMT closed out the pop up window and restarted instruments program.  RTS 1432 GMT.

2/11/16

JEM

The new air compressor for the dry air system has been received and needs to be installed.  Installed the new air compressor from 2100 GMT to 0030 GMT on 2/12/16 GMT.  Leak checked the system and verified OK.  System was working as expected and RH values were low.

2/2/16

KT

The Air compressor belt had broken which provides air flow to the instrument keeping the RH value down. At 1630 GMT there were no replacement belts available to repair the compressor. James Martin and Matt Gibson located a spare air compressor that was installed while a new main Air Compressor could be ordered. As of now the instrument has air flow and is keeping the RH value down to normal values. RTS at 1645 GMT.

2/1/16

KT

Instrument mentor stated that the RH value for the instrument was steadily on the rise. At 2105 GMT found the belt for the compressor that keeps air flowing through the instrument had broken. Troubleshooting still in progress to locate a replacement belt for the air compressor. RTS at 2130 GMT.

2/1/16

KT

Instrument Mentor reported puzzling measurements coming from the instrument and requested help to perform testing. At 2011 GMT placed a 2"x4" aluminum foil square to cover the diffuser. At 2021 GMT the aluminum foil was removed. RTS at 2026 GMT.

12/31/15

JEM

The Central Facility lost power due to an ice storm.  The SAS-Ze did not resume normal operation after power was restored.  A Saspect error was observed on the screen.  Rebooted the SAS-Ze computer @1414 GMT.  At 1415 GMT the SAS-Ze status was observed normal and data collection had resumed.  RTS @1415 GMT.

12/15/15

MS

Computer needed to be restarted following scheduled software and security patch updates.  At 1404 GMT computers were updated and rebooted by the IT department, some data might be lost.

11/17/15

MS

A 7 hour power outage starting at 0700 GMT and power was returned at 1435 GMT, 11/17/2015. Instrument was down. During that time some data might be lost?  Started program back up and RTS 1533 GMT, 11/17/2015.

9/23/15

KT

The computer program was not running and just shut down. At 1235 GMT restarted the computer program. At 1245 GMT the program started and the instrument was online and operational. RTS at 1250 GMT.

9/11/15

CAM

No data had been collected for 3 hours.  Logged in remotely and found SASpect program had not updated since 1539 GMT and displayed the following error: “Unhandled exception has occurred in your application.  If you click Continue, the application will ignore this error and attempt to continue.  If you click Quit, the application will close immediately.  Specified minimum value is equal to the maximum value.  Parameter name: minimum”.  Selected Continue and error closed, but program still did not update.  Selected Run Continuous in SASpect program but same error occurred immediately.  Selected Quit and restarted SASpect.  Program resumed normal operation at 2010 GMT.

8/14/15

CAM

No data collected from SAS-Ze since 1809 GMT.  Logged in remotely and found SASpect program had not updated since 1635 GMT and displayed the following error: “Unhandled exception has occurred in your application.  If you click Continue, the application will ignore this error and attempt to continue.  If you click Quit, the application will close immediately.  Specified minimum value is equal to the maximum value.  Parameter name: minimum”.  Attempted to restart SASpect program at 1954 GMT but another exception occurred.  Rebooted computer at 1955 GMT.  Program started normally on restart.  RTS at 1957 GMT.

8/10/15

CAM

No data had been collected from SAS-Ze since 1609 GMT on 8/8.  Remote status check found SASpect program was not updating and the following error message was displayed: “NationalInstruments.DAQmx.DaqException: Attempted to read sample values that are no longer available.  The requested sample was previously available but has since been overwritten”.  Closed error message and program resumed normal operations at 1332 GMT.

8/3/15

CAM

No data had been collected from SAS-Ze for two hours.  Remote status check found SASpect program had not updated since 1118 GMT and the following error message was displayed: “NationalInstruments.DAQmx.DaqException: Attempted to read sample values that are no longer available”.  Closed error message and program resumed normal operations at 1507 GMT.

7/28/15

CAM

No data was being collected from SAS-Ze.  Remote check using RAdmin found SASpect program was not updating.  Last time shown was 0504 GMT on 7/25.  Closed SASpect program and restarted.  RTS status O.K. at 1402 GMT.

7/27/2015

MS

Program and graphs were not running or updating at 1700 GMT 7/25/2015.  Answered a pop up window and program and graphs started back up at 1317 GMT 7/27/2015.

7/6/15

JEM

The SAS-Ze was red on DSView and was observed with the following error message on the laptop screen.  The specified operation cannot be performed because a task has been aborted or a device has been removed from the system.  There was a power outage overnight that most likely caused this.  Closed the error message and restarted the laptop.  Restarted the SASpect software 1310 GMT.  All data appeared reasonable and the system was RTS @1315 GMT.

6/29/2015

MS

Program not running and spectrum not updating at 1900 GMT 6/27/2015.  Started program back up at 1323 GMT 6/29/2015,  RTS 1323 GMT.

6/22/2015

MS

SAS-Ze program was not on as well as well with the graphs at 0700 GMT 6/19/2015.  Restarted the SAS-Ze program at 1308 GMT on 6/22/2015.  RTS at 1322 GMT.

6/10/15

CAM

Power to all ground mounted transformers is being cut in order to connect the new power panel near the AOS trailer and run power to new PGS building.  Instrument needs to be shut down.  Shut SAS-Ze down at 1351 GMT.  Restarted once power was restored and RTS at 1544 GMT.

6/8/15

MS

SAS-Ze program had stopped and was not collecting data at 0400 GMT on 6/6/2015.  Restarted SAS-Ze Version program at 1303 GMT.  RTS at 1318 GMT.

6/2/15

JEM

The air compressor for the dry air system has started tripping the breaker again.  The mentor is currently on site working on upgrades to the instrument so the compressor problem is not effecting any data.  Troubleshooting determined that the compressor, regulator and start switch were not the cause of the problem and that the new motor is failing.  R/R the compressor motor and RTS the dry air system at 2000 GMT.

6/2/15

MS

SAS-Ze is shut down.    Mentor is on site and troubleshooting SAS-Ze.

5/29/15

CAM

Chiller RH is running high for both SAS-He and Sas-ZE.  T/S found the breaker was tripped for the dry air system compressor.  The breaker was also found tripped yesterday.  Reset breaker and compressor kicked on normally.  Let tank pressurize to 150psi and then manually drained tank to 100psi to force compressor back on. Compressor again operated normally.  Repeated this once more with no problem.  RTS at 1610 GMT.  Will continue to monitor status.

1/23/15

JEM

The air compressor was found inoperable during morning rounds.  R/R the air compressor motor with a newly purchased motor.  RH values began to drop after the compressor was restarted.  RTS @1900 GMT.

1/23/15

KT

The Purge Air Pressure was reading 0. The main outside Air Compressor was off. At 1512 GMT checked the breaker. Breaker didn't appear to be tripped but reset the breaker. No change in status, the Air Compressor did not come on. Pressed the reset switch on the compressor and reset the breaker again. No response from the Air Compressor. Trouble shooting in progress. RTS at 1530 GMT.

1/12/15

JEM

Ken Teske notified me that on morning rounds he found the air compressor inoperable again.  Troubleshooting found the start capacitor had failed again.  Due to continued issues with the air compressor motor, I ordered 2 new motors, one for use and one for a spare.  Was able to find another capacitor that was close enough to the correct size to get the motor running again.  Will install the new motor either when it arrives or when the current motor fails again, whichever happens last.  RTS @2000 GMT.

1/12/15

KT

The main air compressor that provides air for the Dry Air system was not operating. During inspection the breaker for the main air compressor was tripped. Attempted to reset the breaker twice and still could not get the air compressor to return to service. Trouble shooting is in progress. RTS at 1717 GMT.

12/9/14

JEM

The Chiller RH began climbing again approximately 1 hour after the air compressor was returned to service on 12/8.  Tried resetting the breaker and again it would trip after a few seconds of operation.  Disassembled the motor and checked the internal centrifugal start switch again and it was not stuck and was working properly.  Suspect a problem with the motor.  R/R the motor with the spare.  Reset the breaker after installing and connecting the spare motor.  The breaker would immediately and the motor would not even try to run.  Removed the motor and took both motors to the repair lab for further inspection.  The spare motor would also trip the breaker in the repair lab and further inspection found a burnt spot in one of the windings causing a short.  Powered on the original motor in the lab and it ran just fine and current draw was only 7 amps.  Returned the original motor to the optical trailer air compressor and again it would trip the breaker after a few seconds of operation.  Current draw was measured at ~40 amps.  Inspection of the cut in/out switch connected to the high pressure cutoff revealed that all of the contact pads had a large amount of carbon buildup and damage from arcing.  R/R the cut in/out switch with a new one.  Due to all of the times the breaker has tripped in the past the breaker was also replaced with a new one.  Powered the compressor on at 2200 and it ran normally.  Dry air flow returned to the SAS-Ze and the RH began to drop.  RTS @2200 GMT.

12/8/14

JEM

Received an email from the mentor that the RH in the SASZHe Chiller began to climb on 12/5 and was currently very high.  Checked the air compress for the dry air system and it was not operating.  Tried resetting the breaker several times and the breaker would trip after a few seconds of operation.  Disassembled the air compressor motor and the internal centrifugal switch was stuck in the on position, not allowing the start winding to disengage once the motor was up to speed causing excessive current draw.  Replaced the internal centrifugal start switch with a new one.  Reassembled the motor and powered on the compressor.  Dry air flow returned to the SASZHe and the RH began to drop.  RTS @1830 GMT.

12/1/14

JM/CM

The dry air compressor has tripped the breaker several times in the past few weeks.  Replaced the start capacitor for the compressor motor at 1900 GMT.  One of the terminals broke off the original capacitor when it was disconnected, so it may have been the source of the problem.

12/1/14

KT

The Outside Air Compressor was not running. At 1520 GMT found that breaker had been tripped. Reset the breaker and the Air Compressor powered back on. RTS at 1535 GMT.

11/12/14

JEM

Was notified by site ops that the air compressor used to supply dry air to the freezer is not working.  The circuit breaker was found tripped and when trying to reset the breaker it trips again after about 1 second.  Troubleshooting found that the start capacitor on the motor has failed.  R/R the start capacitor and reset the circuit breaker.  Compressor began working normally.  After the compressor pressured up, observed the freezer humidity begin to drop.  RTS @1830 GMT.

10/1/14

CAM

The dry air system compressor motor failed on 9/30.  Removed the failed motor and replaced with a spare rebuilt motor.  Removed smaller compressor that had been temporarily plumbed in yesterday and placed the original compressor back in service at 2000 GMT.

9/30/14

CAM

Technician found during daily rounds that the compressor motor for the dry air system had failed.  Temporarily connected a smaller compressor to the dry air system until the large compressor can be repaired.  RTS at 2115 GMT.

8/14/14

CAM

The RH inside the SAS-He refrigerator has risen over the past several weeks and is at times near 60%.  Found that SAS-He, SAS-Ze, SORTI, and SWS were all on the same regulated dry air line.  There was another regulated air line that had only RSS on it.  Reconfigured the air lines so that SAS-He and SORT were on a line by themselves.  RH dropped to near 0% for both SAS instruments and appears to be holding.  RTS at 2130 GMT.

7/25/14

KT

Instrument mentor sent out a sunshade baffle mount to be installed. At 1306 GMT removed the old baffle mount and installed new baffle mount. RTS at 1326 GMT.

5/28/14

Mentor

Instrument mentor on site and reinstalled SAS-Ze.

5/2/14

Mentor

Flynn

The instrument needs to be sent to the vendor for calibration. Powered off and removed the instrument for calibration at 1400 GMT.

4/21/14

CAM

Instrument mentor requested information on the USB/RS232 and RS232/RS422 converters in the SAS instruments.  During inspection, found that the ground wire for the SAS-Ze motion control motor was disconnected.  This motor is not in use, but the ground should be connected.  Stopped program and shut system down at 1957 GMT.  Secured ground connection, restored power and RTS at 2006.

3/4/14

KT

System shut down due to power outage that occurred on 3/2/2014 at 2120 GMT. CF was closed 3/3/2014 due to snow storm. CF opened on 3/4/2014. Powered on the computer at 1351 GMT. System came online at 1354 GMT. RTS at 1406 GMT.

1/30/14

CAM

Kay Electric needed to shut the power off briefly while they worked on the meter.  Power was shut off at 1614 GMT and restored at 1623 GMT.  The optical trailer UPS died during that time, causing instruments to lose power.  Power cycled UPS at 1647 GMT to restore power.  SAS-Ze restarted and RTS at 1651 GMT. 

1/22/14

Mentor

Instrument mentor (Flynn) on site for 2 days to reinstall instrument.  Instrument was installed on the east optical trailer platform next to the SWS.  RTS at 1700 GMT.

7/2/12

CAM/KT

Instrument mentor requested the SAS-Ze be removed and sent to PNNL, where it will be modified for AMF2 deployment on the MAGIC campaign.  Removed instrument at 1500 GMT and delivered to S/R to be sent to PNNL.

2/13/12

PAD/KT

At 1542 GMT, found SAS-Ze computer was down.  Troubleshooting found laptop power supply had failed.  Installed temporary power supply provided by computer dept.  RTS at 1622 GMT.

1/17/12

PAD

Mentor requests refrigerator door replaced with modified door supplied by mentor.  Modified door allows dry air to flow into refrigerator.  Stopped program and shut down system at 1935 GMT.  Installed door.  Started system but instrument was not in correct position to shade diffuser.  Az stall and Band 1 High status boxes are red after restart.  Stopped/ restarted program with same results.  Reset azimuth offset.  Instrument is now in correct alignment and shading correctly.  RTS at 2035 GMT.

1/4/12

PAD

Band 1 low and high boxes are red.  Instrument is not in correct azimuth alignment to shade window.  Stopped program and zeroed azimuth.  Initialized and reset azimuth offset.  Started measurements and shadowband is now shading correctly.  RTS at 1547 GMT.

1/3/12

PAD

Error shown:  System Argument out of Range Exception.  Cleared error and program finished taking measurement.  Rebooted computer but azimuth is not correct.  Attempted to reset azimuth offset but it would not go to zero when initialized.  Able to use software to move azimuth both directions but it still will not initialize.  Entered last known offset for azimuth and started measurement.  Azimuth position is still not correct and Az stall and Band 1 High 1.8 boxes are red.  Notified instrument mentor.  RTS at 1626 GMT.

12/29/11

PAD

Band 1 low and high status boxes are red.  Stopped program and zeroed inner band.  Started program and shadowband is shading correctly but Band 1 low status box is red and shows 187 deg.  RTS at 1545 GMT.

12/28/11

PAD

Error -3 is shown on monitor.  Spectrometer error and band 1 low and band 1 high status boxes are red.  Cleared error and stopped program.  Rebooted computer.  After starting measurements, shadowband was not in correct position.  Stopped measurements and zeroed shadowband.  Started measurements and shadowband is now shading correctly.  RTS at 1607 GMT.

12/21/11

PAD

Shadowband is horizontal and not shading correctly.  Band 1 low and high status boxes are red.  Zeroed inner band and it stopped in correct vertical position but showed -90.5 deg.  Changed inner band offset from -90.5 to 1.49.  Restarted program and azimuth stall error was shown.  Stopped program and zeroed azimuth and inner band.  Started program and shadowband and azimuth are now correct and window is shaded.  Instrument took one series of measurements, then shadowband moved to horizontal and stayed.  At 1603 GMT, observed shadowband not shading.  Zeroed inner band and started program.  Instrument again moved to correct azimuth and shadowband position and took one series of measurements.  Shadowband then moved to horizontal and stayed even though program indicates shadowband movement.  Notified mentor via email.  Waiting on reply.  RTS at 1623 GMT.

12/1/11

PAD

Shadowband is not in correct position.  Band 1 Low and Band 1 High status boxes are red.  Stopped/ restarted SASpect program.  No change.  Rebooted computer and after program started, shadowband moved to horizontal and stayed there.  Shut down system completely and restarted.  Az stall and Band 1 High boxes are red.  Consulted with mentor.  Zeroed azimuth and shadowband.  Restarted program and instrument moved to correct azimuth position and shadowband is now shading.  RTS at 1716 GMT.

10/27/11

PAD

Overcast conditions and unable to verify shading but instrument azimuth position is clearly not correct.  Motion tab shows azimuth position of 28.669.  Stopped program and rebooted computer.  After restart, azimuth position of 29.918 is shown but instrument appears to be in correct position.  RTS at 1426 GMT.

10/26/11

PAD

Last SDS collection was at 1300 GMT.  At 1512 GMT, instrument had power but computer was down.  TS found computer power supply had failed.  Borrowed SAS-He power supply since that system in not in service.  Notified instrument mentor.  Removed Dell Model DP/N OCM899, SN CN-OCM899-73245-06A-1564-A01.  Installed Dell Model DP/N OCM899, SN CN-OCM899-73245-069-5562-A01.  RTS at 1527 GMT.

10/24/11

PAD

Motor error, Band 1 low, and Band 1 high status boxes are red.  Status shows "Acquire D blocked" but shadowband is not shading window.  Stopped SASpect program and rebooted computer.  Program restarted correctly and shadowband is now shading.  RTS at 1542 GMT.

9/27/11

PAD

Error -3 shown.  Status shows:  Running Dark f.  Acquiring 12 of 60.  Cleared error but it keeps coming back.  Used task manager to stop SASpect program.  Rebooted computer.  Program started correctly and shadowband is shading.  RTS at 1554 GMT.

9/26/11

PAD

Status shows Acquiring 21 of 150, Acquire D blocked.  Shadowband is on wrong side of instrument to shade sun.  Stopped SASpect program and rebooted computer.  Instrument is not shading correctly.  RTS at 1406 GMT.

9/20/11

PAD

System status shows "Acquire D blocked", but shadowband is horizontal and not shading window.  Stopped SASpect program and rebooted computer.  Program started correctly and shadowband is now shading.  RTS at 1454 GMT.

9/19/11

PAD

SASpect program show instrument is operating but shadowband is not shading and does not reflect position shown on program.  Tried to stop program but after it finished current measurements error was repeatedly shown:  Cannot write to closed text file.  Used task manager to close SASpect program.  Rebooted computer and program started correctly.  Shadowband is now shading correctly.  RTS at 1550 GMT.

9/14/11

PAD

Shadowband is not shading correctly.  Stopped program and rebooted computer.  Program started and shading is now correct.  RTS at 1552 GMT.

9/7/11

PAD

Error shown:  System argument out of range exception.  Index out of range.  Must be non-negative and less than the size of the collection...  Took screenshot of error message and saved on desktop.  Cleared error and rebooted computer.  Program started correctly.  RTS at 1424 GMT.

8/15/11

PAD

Program show Acquire D blocked.  Shadowband is horizontal and not shading window.  Stopped/restarted Saspect program but shadowband is still horizontal.  At 1412 GMT, error shown:  Azimuth hard limit not found.  Stopped program and rebooted computer.  Program did not startup on computer reboot.  Rebooted computer again and Saspect program started correctly.  Shadowband is now shading window.  RTS at 1427 GMT.

8/12/11

PAD

Shadowband is horizontal at 89.9 degrees and is not shading window.  Status shows Acquire D blocked, 81 of 150.  Stopped/restarted Saspect program.  Shadowband moved into correct position and is now shading window.  RTS at 1336 GMT.

8/9/11

PAD

CF lost power on 8/8/11 at 2247 GMT.  Power was restored on 8/9/11 at 0211 GMT.  Found computer inactive.  At 1522 GMT, powered up computer.  Program started correctly and shadowband is shading.  RTS at 1537 GMT.

8/2/11

PAD

Shadowband is horizontal and not blocking sunlight.  Stopped experiment and re-zeroed Azimuth and Inner Band.  Started measurement, inner band is now in correct position, but Azimuth is not correct.  Notified mentor at 1540 GMT.  Mentor remotely accessed computer and determined program had the wrong Azimuth offset.  Mentor found correct offset in previous Saspect program file and entered it into current program.  Azimuth position is now correct.  RTS at 1620 GMT.

7/5/11

PAD

SASpect program is shown but not running.  Error shown: Error-3.  Cleared error and another error was shown:  Argument was out of range, must be non-negative and less than size of collection.  Cleared errors but they both keep popping up repeatedly.  Rebooted computer and now Error-4 is shown.  At 1755 GMT, remotely accessed  computer and rebooted.  Program started correctly.  RTS at 1810 GMT.

6/27/11

PAD

SDS shows last successful collection was 2 days ago.  Remotely accessed computer and error was shown:  System argument out of Range Exception.  Index out of range.  Must be non-negative and less than the size of collection.  Cleared error and program started taking measurements.  RTS at 1405 GMT.

5/16/11

PAD

SAS-Ze azimuth drive has failed.  Mentor has designed and built a baffle to prevent direct sunlight from entering window and requests it be installed.  Installed baffle that was sent from Mentor as instructed at 1932 GMT.  RTS at 1947 GMT.

5/3/11

PAD

Instrument shadowband optical switches have not been working.  Removed cover and tested voltage at optical limit switch.  Voltage measured .732 VDC at the switch connector instead of 5VDC expected.  Notified mentor.  At 1754 further troubleshooting found bad wire between azimuth drive housing and head.  Broken or bad wire was bypassed by lab technician.  Put all covers in place and started instrument at 2024 GMT with mentor remotely accessing computer.  Shadowband is now working correctly but azimuth drive is not responding.  Drive will move instrument a few degrees and then sounds like something is slipping or binding.  Drive will not always turn in intended direction.  Placed all covers back on instrument.  Troubleshooting at this time.  RTS at 2144 GMT.

4/12/11

PAD

At 1505 GMT, ratcheting sound was observed.  Slight pressure to Azimuth drive would change sound.  Stopped experiment and left voicemail with mentor.  Mentor later called and we reset the zero for the Azimuth.  While resetting inner band zero, band would proceed ~20 degrees past horizontal and try to continue.  Removed side of instrument enclosure and found machined limit plate was loose on shaft.  Tightened limit plate to shaft on flat side.  Tried to zero inner band but it would proceed to horizontal and then try to continue.  Installed panel on instrument enclosure.  Troubleshooting with mentor at this time.  RTS at 1735 GMT.

3/11/11

Mentor

Instrument installed by Connor Flynn and Dan Nelson, 3/9-3/11.